OpenStudy (anonymous):

Explain how to get the product of 2 times the square root of 5 and 4 times the square root of 10.

OpenStudy (anonymous):

\[2\sqrt{5} \times 4\sqrt{10}\]\[8\sqrt{50} = 8\sqrt{2 \times 25} = 8 \times 5 \times \sqrt{2} = 40\sqrt{2} \approx 56.56\]